### v1

# CIF-Bench: A Chinese Instruction-Following Benchmark for Evaluating the Generalizability of Large Language Models

Abstract
The advancement of large language models (LLMs) has enhanced the ability to generalize across a wide range of unseen natural language processing (NLP) tasks through instruction-following. Yet, their effectiveness often diminishes in low-resource languages like Chinese, exacerbated by biased evaluations from data leakage, casting doubt on their true generalizability to new linguistic territories. In response, we introduce the Chinese Instruction-Following Benchmark (CIF-Bench), designed to evaluate the zero-shot generalizability of LLMs to the Chinese language. CIF-Bench comprises 150 tasks and 15,000 input-output pairs, developed by native speakers to test complex reasoning and Chinese cultural nuances across 20 categories. To mitigate evaluation bias, we release only half of the dataset publicly, with the remainder kept private, and introduce diversified instructions to minimize score variance, totaling 45,000 data instances. Our evaluation of $28$ selected LLMs reveals a noticeable performance gap, with the best model scoring only 52.9%, highlighting the limitations of LLMs in less familiar language and task contexts. This work aims to uncover the current limitations of LLMs in handling Chinese tasks, pushing towards the development of more culturally informed and linguistically diverse models with the released data and benchmark [^1] https://yizhilll.github.io/CIF-Bench/.
Figure 1: A large language model can tackle English task translated to Chinese, but fail to respond to instruction originally in Chinese.
1 Introduction
The landscape of natural language processing (NLP) has been dramatically reshaped by the emergence of large language models (LLMs), which have demonstrated an ability to generalize across unseen NLP tasks, often showcased through the framework of instruction-following Mishra et al. (2021); Sanh et al. (2021); Wei et al. (2021). Despite these advances, skepticism remains regarding the transferability of this instruction-following capability, particularly in multilingual contexts. The models perform worse when switching to Chinese due to the prevalence of English training data Huang et al. (2023b); Zhang et al. (2023b), as figured in Fig. 1. This concern is exacerbated by observations that benchmarks designed to assess the capabilities of LLMs may inadvertently suffer from biased evaluations due to data leakage Sainz et al. (2023), particularly when web-scale datasets are employed to enhance model generalizability Raffel et al. (2023). Such observations raise a critical question: While the generalizability of LLMs appears intriguing, do these models face significant challenges when evaluated in less-common language contexts with newly curated instruction-formatted tasks?
To answer this question, we introduce the Chinese Instruction-Following Benchmark (CIF-Bench), a novel benchmark designed for the zero-shot generalizability evaluation of LLMs, with Chinese serving as an insightful example for multilingual transferred instruction-following tasks. Our benchmark comprises $150$ tasks and $15,000$ input-output pairs, with the assistance of native speaker annotators, ensuring the inclusion of human-authored tasks that are not only challenging but also naturally expressed. A significant portion ($38.7\%$) of these tasks are designed to test a model’s complex natural language inference (NLI) and reasoning capabilities, as well as drawing upon Chinese culture spread across 20 distinct categories. In an effort to mitigate future evaluation biases from data leakage, we decide to publicly release only half of the data instances, reserving the rest as a private dataset to maintain an impartial benchmark. Furthermore, CIF-Bench enhances its robustness by introducing 5 variations of instructions per task, using these to diminish score variance in private split evaluations as discussed in §5. CIF-Bench also pioneers a model-based automatic pipeline designed to tackle the inherent challenges of evaluating open-ended natural language generation outputs Gehrmann et al. (2021).
By selecting a range of popular LLMs that support Chinese for evaluation, we aim to depict the limits of current instruction-following capabilities in language transfer contexts as the many models follow an English-oriented pre-training paradigm Huang et al. (2023b). Our findings reveal that even the best-performing model achieves a score of only $52.9\%$ on CIF-Bench, underscoring the gap that exists when LLMs are confronted with tasks in a less-familiar language and unseen data instances. We find that this performance decrement is particularly noticeable in scenarios involving unseen tasks and unseen input-output pairs, contrasting with the models’ performance on existing Chinese datasets and translated English-language tasks. Such results suggest that while LLMs exhibit impressive generalizability in a context more aligned with observed data, their effectiveness diminishes when faced with the dual challenges of unacquainted languages and novel tasks.
To summarize our contributions, we:
  • Present a new benchmark that addresses a critical gap in existing NLP research by focusing on the generalizability of LLMs to an underrepresented language in terms of training and evaluation resources;
  • Construct an instruction-following evaluation dataset with $150$ tasks and $45,000$ data samples, and release half of the input-output pairs for future LLM evaluation research;
  • Provide an in-depth analysis of 28 LLMs, revealing their limitations in adapting to less familiar languages and task contexts, offering insights into where improvements are needed for instruction-following generalizability.
2 Related Work
Instruction-Following Evaluation.
Large-scale pre-trained language models have been found that they can generalize across unseen tasks by fine-tuned on formatted task instructions (Khashabi et al., 2020; Mishra et al., 2021; Wei et al., 2021; Sanh et al., 2021). Early studies attempt to fine-tune and evaluate such a capability in a few-shot manner by providing input-output examples Ye et al. (2021); Mishra et al. (2021). Following that, another line of research Bach et al. (2022); Wang et al. (2022b) improves the evaluation reliability from the perspective of scaling the task quantity and providing well-defined corresponding instructions. A more recent concurrent work FollowBench Jiang et al. (2023) proposes to use a similar that crafting multiple instructions for a task to evaluate the LLMs. A core distinction between CIF-Bench and this strategy is that whilst the FollowBench aims at applying different kinds of constraints in the same task, we focus on assessing whether models can stably perform given diversely expressed instructions with the same semantics.
Chinese LLM Benchmarks.
There have been important efforts, such as CLUE (Xu et al., 2020) and CUGE (Yao et al., 2021), made to evaluate the pre-trained language on extensive tasks in the Chinese context, which consider the traditional taxonomy of natural language understanding and generation. As these benchmarks are restricted in the prediction formats and could not fully measure the cross-task generalization of LLMs in the free-form outputs, more recent studies Huang et al. (2023b); Li et al. (2023) propose to reformat the tasks into multi-choice question answering, mostly examining the knowledge-base abilities in Chinese. However, such a strict format could impede the models from fully generalizing to more complex reasoning and creative tasks. Thereby, we argue that there is a lag in evaluating LLMs instruction-following capacity in the Chinese language.
3 The Challenging Chinese Instruction-Following Benchmark
The Challenging Chinese Instruction-Following Benchmark unifies the NLP tasks in the prompt-based instruction-following schema Mishra et al. (2021) and evaluates the LLMs in a zero-shot manner, which is to say that the models are expected to directly provide the correct output given the concatenation of the task instruction and data input texts. Formally, for each data sample in CIF-Bench, the three components we refer to are:
  • An instruction that is provided as the introductory information for a specific NLP task, which is an implicit definition of a “mapping function" (i.e., task background context) that must be interpreted by the models before proceeding.
  • An span of input text that encompasses the context to define the specific task scenario.
  • A reference as the (potentially) standard output in the data instance.
Table 1: The Statistics of CIF-Bench instruction data. #Instruction and #Input-Output refer to the quantity of examples contained in each task.
Splitnormal-$\rightarrow$ Private Public
#Task 150
#Instruction 5 1
#Input-Output 50 50
Total Instances 37,500 7,500
We define a total of 150 curated tasks, constructed according to Chinese linguistic and societal backgrounds, as well as from existing NLP tasks in Chinese and English. To improve the evaluation robustness, we provide a diversified set of $5$ instructions with the same semantics for each task. Considering the potential data leakage issue of LLM benchmarks, we split two halves of $100$ input-output pairs in each task into private and public partitions, and only test and release the public split which contains one instruction variant. In sum, there are $45,000$ human-annotated [instruction, input, output] instances produced in CIF-Bench, as suggested in Table 1. In addition, we provide detailed instructions for all the tasks in § A.1.
Table 2: The statistics of existing and newly designed Tasks. The existing tasks and instances include those translated from English as well as original Chinese data.
Task Instance
Existing Newly Annotated
Existing (113) 5,650 5,650
Newly Designed (37) N/A 3,700
Total 5,650 9,350
3.1 Data Collection
Annotation Protocol
CIF-Bench is designed for the extensive evaluation of Chinese comprehension and generation capabilities in LLMs, particularly focusing on aspects such as creative generation and linguistic abilities that existing benchmarks, such as C-Eval Huang et al. (2023b) and C-MMLU Li et al. (2023), struggle to assess. First, we select 113 diverse existing English NLP tasks, as shown in Table 1 from Super Natural Instructions (SNIWang et al. (2022b) and other research work (full list in § A.1). We then describe these task instructions in Chinese and a semantically balanced distributed subset from each original English NLP task as the Public split of CIF-Bench. We further ask expert native Chinese speakers, who minimally have undergraduate degrees, to annotate 100 samples per task based on the translated task instructions. These samples are further deduplicated according to their semantic embeddings. We finally select 50 samples per task as the Private split of CIF-Bench, to guarantee each sample’s validity and the balanced label distribution of each task.
To further improve CIF-Bench’s task diversity, we create 37 additional new tasks and state the related Chinese instructions. Specifically, we focus on adding Chinese tasks about Creative Natural Language Generation, Traditional Chinese, and Complex Role-Playing Text Games. We ask the expert native speakers to annotate 200 samples per task based on the translated task instructions. These samples are deduplicated and we further select the Public and Private split from it. Each task is further annotated with 4 Private paraphrased instructions to test whether LLMs understand the Chinese instructions’ meanings or overfit to the instructions in the Public split. Each sample and instruction is manually guaranteed or written by at least one of the authors to make sure that CIF-Bench is reliable.
Figure 2: Task Category Distribution in CIF-Bench. The radii have three groups, determined by the number of tasks contained ($\leq 10$, $\leq 20$, and $>20$).
3.2 Task Category
Whilst diverse tasks are provided in CIF-Bench, it would be difficult to analyze the extensive scores from all of the tasks. By reviewing and summarizing the existing NLP tasks and instruction-following benchmarks, we accordingly categorize the $150$ tasks into 20 basic types in a multi-label fashion (i.e., a task can be belong to more than one category). Each category consists of $2$ to $36$ tasks and the quantity distribution is revealed in Figure 2. Other than the $36$ “commonsense” tasks requiring a wide-ranging knowledge base, there are two dominant categories that aim to challenge the logical reasoning abilities of LLMs in CIF-Bench, including $30$ ”natural language inference (NLI)” and $29$ “reasoning” tasks. In particular, there are 18 tasks designed to require knowledge of unique Chinese cultural contexts. We describe the definition of each category and the task numbers in § A.2.
Figure 3: An Exemplar Prompt for GPT-4 Evaluator for the Task “Chinese Rhetoric Detection”.
Table 3: Overall results in CIF-Bench Private split with diversified instructions (1/2). The first column is the average score across all the tasks, and the other columns are average scores grouped by task categories. The cells are highlighted with fading colors from maximum to minimum in a column.
Model Name Overall Chinese Culture Classification Code Commonsense Creative NLG Evaluation Grammar Linguistic Motion Detection NER
Baichuan2-13B-Chat 0.529 0.520 0.674 0.333 0.641 0.497 0.686 0.542 0.528 0.578 0.563
Qwen-72B-Chat 0.519 0.486 0.630 0.296 0.634 0.508 0.634 0.458 0.520 0.494 0.550
Yi-34B-Chat 0.512 0.483 0.606 0.347 0.623 0.497 0.598 0.480 0.490 0.575 0.525
Qwen-14B-Chat 0.500 0.481 0.582 0.307 0.614 0.494 0.645 0.428 0.475 0.496 0.513
Deepseek-LLM-67B-Chat 0.471 0.467 0.571 0.259 0.577 0.486 0.549 0.442 0.476 0.475 0.509
Baichuan-13B-Chat 0.450 0.408 0.491 0.286 0.552 0.439 0.670 0.417 0.422 0.482 0.486
Chatglm3-6B 0.436 0.381 0.439 0.330 0.541 0.452 0.577 0.310 0.358 0.436 0.453
Yi-6B-Chat 0.417 0.402 0.454 0.313 0.523 0.425 0.506 0.383 0.383 0.487 0.396
Baichuan2-7B-Chat 0.412 0.437 0.647 0.160 0.520 0.402 0.580 0.511 0.444 0.455 0.407
Chatglm2-6B 0.352 0.278 0.469 0.346 0.403 0.424 0.535 0.274 0.397 0.406 0.240
Chatglm-6B-Sft 0.349 0.265 0.454 0.365 0.385 0.462 0.554 0.296 0.379 0.427 0.232
Chinese-Llama2-Linly-13B 0.344 0.250 0.462 0.311 0.399 0.429 0.557 0.273 0.358 0.385 0.268
GPT-3.5-Turbo-Sft 0.343 0.269 0.427 0.298 0.389 0.395 0.575 0.325 0.365 0.389 0.226
Chinese-Alpaca-2-13B 0.341 0.242 0.421 0.356 0.382 0.442 0.602 0.256 0.363 0.430 0.210
Chinese-Alpaca-13B 0.334 0.250 0.399 0.348 0.364 0.435 0.616 0.275 0.349 0.421 0.223
Chinese-Alpaca-7B 0.334 0.216 0.412 0.378 0.381 0.425 0.576 0.265 0.359 0.393 0.243
Chinese-Llama2-Linly-7B 0.333 0.218 0.451 0.330 0.396 0.427 0.583 0.248 0.350 0.410 0.231
Tigerbot-13B-Chat 0.331 0.205 0.397 0.309 0.385 0.420 0.614 0.310 0.379 0.341 0.276
Telechat-7B 0.329 0.267 0.338 0.321 0.420 0.404 0.420 0.272 0.265 0.327 0.320
Ziya-Llama-13B 0.329 0.196 0.402 0.324 0.341 0.428 0.616 0.312 0.349 0.400 0.228
Chinese-Alpaca-33B 0.326 0.234 0.370 0.372 0.364 0.429 0.614 0.246 0.318 0.377 0.221
Tigerbot-7B-Chat 0.325 0.218 0.395 0.306 0.370 0.413 0.631 0.294 0.370 0.368 0.215
Chinese-Alpaca-2-7B 0.323 0.215 0.374 0.335 0.366 0.415 0.546 0.257 0.326 0.395 0.215
Aquilachat-7B 0.309 0.162 0.234 0.291 0.320 0.437 0.344 0.135 0.266 0.309 0.287
Moss-Moon-003-Sft 0.302 0.214 0.405 0.274 0.347 0.380 0.448 0.305 0.341 0.378 0.232
Qwen-7B-Chat 0.301 0.211 0.410 0.289 0.349 0.391 0.531 0.219 0.387 0.404 0.208
Belle-13B-Sft 0.264 0.198 0.307 0.285 0.316 0.349 0.409 0.237 0.305 0.222 0.177
CPM-Bee-10B 0.244 0.234 0.377 0.024 0.278 0.311 0.255 0.302 0.278 0.327 0.148
3.3 Task-based Automatic Evaluation
As the CIF-Bench aims to provide a comprehensive evaluation of the LLM instruction-following capability, we argue that the metrics should be designed case by case in task granularity to evaluate the open-ended textual outputs, rather than simply reformatting all tasks into choice questions and using the conditional probability to approximate the models’ predictions.
After a thorough review of the task instructions, we categorize the output requirements into the four following types and design corresponding task-level metrics. Multi-class Classification: We use accuracy as the metric if the task requires the model to predict one label from 2 or more classes in the output. Multi-label Classification: We use F1 score as the metric if the task requires the model to predict one label from 2 or more classes in the output. Creative Generation: Regarding the tasks that have no absolute criteria of the standard answer, we require a model-based evaluator to provide information regarding a given output, including creativity, fluency, the level of instruction-following, and the confidence of the evaluator. Semantic Similarity: For the remaining tasks that can be evaluated by the semantic similarity between the golden reference and model output, we use a pre-trained language All scores used in CIF-Bench either naturally range from 00 to $1$, or are normalized to the same range.
One core dilemma in evaluating the open-ended instruction-following capabilities of LLMs is that model predictions are hard to verify even with reference answers. For instance, it is intractable to handcraft regex rules to extract the predictions from LLMs for the extensive number of tasks, since the answers could be expressed in various formats, or drowned in redundant contexts like reasoning progress. Inspired by G-Eval Liu et al. (2023), we leverage OpenAI’s GPT-4222https://openai.com/gpt-4 as a relatively reliable evaluator for multi-class classification, multi-label classification, and creative generation tasks, to overcome such issues. The GPT-4 evaluator is prompted to assess the outputs according to the given task instruction and the input-output reference, as shown by the example in  Figure 3 and the full list of evaluation prompts in § A.1. For the answers that can be evaluated with semantic similarity, we use a lightweight multilingual encoder, BLEURT Sellam et al. (2020), to measure the relevance between the reference and LLM output.
Given a set of task instructions $I$, we denote the performance score of model $m$ on task $t$ as:
$S^{m}_{t}=\frac{1}{|D_{t}|}\sum_{d\in D_{t}}\frac{1}{|I|}\sum_{i\in I}{{s^{m}_{t}(i,d)}}$
, where $D_{t}$ refers to the set of data samples for task $t$. In the case of the public split, the instruction set $I$ is reduced to one single element. In we take the average of task-level scores $\overline{S^{m}}$ as the indicator of overall performance for a model $m$.
Table 4: Overall results in CIF-Bench Private split with diversified Instructions (2/2). The first column is the average score across all the tasks, and the rest columns are average scores grouped by task categories. The cells are highlighted with fading colors from maximum to minimum in a column.
Model Name Overall NLI QA Reasoning Role Playing Sentiment Structured Data Style Transfer Summarization Toxic Translation
Baichuan2-13B-Chat 0.529 0.632 0.569 0.515 0.752 0.624 0.459 0.462 0.332 0.441 0.273
Qwen-72B-Chat 0.519 0.626 0.565 0.528 0.762 0.613 0.496 0.459 0.282 0.608 0.271
Yi-34B-Chat 0.512 0.619 0.554 0.494 0.757 0.580 0.472 0.439 0.346 0.514 0.259
Qwen-14B-Chat 0.500 0.616 0.548 0.507 0.764 0.583 0.469 0.453 0.283 0.575 0.262
Deepseek-LLM-67B-Chat 0.471 0.566 0.496 0.439 0.711 0.546 0.409 0.436 0.262 0.570 0.235
Baichuan-13B-Chat 0.450 0.565 0.505 0.377 0.704 0.552 0.387 0.402 0.350 0.431 0.304
Chatglm3-6B 0.436 0.544 0.503 0.414 0.762 0.560 0.446 0.402 0.321 0.391 0.270
Yi-6B-Chat 0.417 0.523 0.457 0.369 0.754 0.482 0.401 0.380 0.310 0.455 0.227
Baichuan2-7B-Chat 0.412 0.489 0.395 0.406 0.670 0.517 0.342 0.298 0.101 0.463 0.138
Chatglm2-6B 0.352 0.397 0.352 0.326 0.714 0.438 0.298 0.313 0.320 0.461 0.190
Chatglm-6B-Sft 0.349 0.380 0.321 0.292 0.718 0.415 0.296 0.333 0.351 0.441 0.190
Chinese-Llama2-Linly-13B 0.344 0.390 0.330 0.313 0.653 0.433 0.279 0.332 0.292 0.457 0.181
GPT-3.5-Turbo-Sft 0.343 0.382 0.394 0.345 0.710 0.433 0.324 0.266 0.290 0.397 0.225
Chinese-Alpaca-2-13B 0.341 0.376 0.334 0.317 0.714 0.459 0.299 0.316 0.308 0.452 0.200
Chinese-Alpaca-13B 0.334 0.370 0.309 0.319 0.724 0.426 0.285 0.307 0.298 0.445 0.181
Chinese-Alpaca-7B 0.334 0.383 0.326 0.295 0.710 0.409 0.301 0.327 0.325 0.405 0.186
Chinese-Llama2-Linly-7B 0.333 0.367 0.345 0.276 0.698 0.433 0.259 0.315 0.310 0.469 0.168
Tigerbot-13B-Chat 0.331 0.363 0.329 0.301 0.694 0.419 0.280 0.310 0.283 0.393 0.186
Telechat-7B 0.329 0.388 0.355 0.244 0.672 0.344 0.334 0.335 0.299 0.364 0.184
Ziya-Llama-13B 0.329 0.351 0.279 0.313 0.721 0.468 0.311 0.291 0.278 0.431 0.175
Chinese-Alpaca-33B 0.326 0.368 0.300 0.314 0.713 0.428 0.288 0.303 0.295 0.401 0.199
Tigerbot-7B-Chat 0.325 0.355 0.313 0.292 0.713 0.415 0.283 0.315 0.290 0.389 0.171
Chinese-Alpaca-2-7B 0.323 0.375 0.318 0.289 0.698 0.417 0.285 0.303 0.312 0.439 0.193
Aquilachat-7B 0.309 0.337 0.342 0.236 0.609 0.255 0.249 0.400 0.527 0.430 0.306
Moss-Moon-003-Sft 0.302 0.317 0.321 0.267 0.694 0.375 0.251 0.259 0.288 0.424 0.152
Qwen-7B-Chat 0.301 0.325 0.297 0.278 0.681 0.419 0.266 0.251 0.248 0.371 0.157
Belle-13B-Sft 0.264 0.317 0.284 0.242 0.631 0.299 0.244 0.222 0.234 0.296 0.133
CPM-Bee-10B 0.244 0.286 0.224 0.147 0.603 0.277 0.117 0.263 0.220 0.352 0.125
4 Experiments
Baselines.
We compare the performance of existing LLMs that have been trained on Chinese corpora. We select ChatGPT, for which we use gpt-3.5-turbo-instruct, [^3] https://openai.com/ which we believe corresponds to instructGPT text-davinci-002. Then we select a series of open-source LLMs, including ChatGLM (Zeng et al., 2023), AquilaChat-7B. [^4] https://github.com/FlagAI-Open/FlagAI/ Baichuan (Baichuan, 2023), Deepseek-Llm-67B-Chat (DeepSeek-AI, 2024), Qwen (Bai et al., 2023), Yi, [^5] https://github.com/OrionStarAI/OrionStar-Yi-34B-Chat/tree/main tigerbot-7b-chat (Chen et al., 2023), TeleChat (Wang et al., 2024), CPM-Bee-10B, [^6] https://github.com/OpenBMB/CPM-Bee, and Moss-Moon (Sun et al., 2023), which have been trained from scratch on a large volume of data in both English and Chinese. We additionally select other instruction-following LLMs, such as Ziya-LLaMA-13B (Wang et al., 2022a), Chinese-Alpaca (Cui et al., 2023), Linly-Chinese-LLaMA2 (Zhao et al., 2023), and BELLE (BELLEGroup, 2023), which are trained with Supervised Fine-Tuning (SFT) on Chinese data, including web texts, books, and code, and then trained via alignment techniques.
Settings.
For inference, we use four Nvidia A100 GPUs with 80GB of VRAM. To optimize GPU resource usage, we directly employed the vLLM framework (Kwon et al., 2023) for LLM inference on CIF-Bench where applicable. This setup enables each model to complete all tasks within approximately 6 to 12 hours. For models not supported by the vLLM, we adhere to the configurations specified in official repositories, resulting in an inference duration ranging from 12 to 48 hours. During the evaluation, we use two Nvidia 2080-Ti 12GB GPUs to conduct the BLEURT semantic similarity calculations, and use the gpt-4-turbo-preview version of GPT-4 API as the open-ended evaluator for the rest of tasks.
Table 5: Comparison between English-translated and newly annotated Chinese tasks in the Public split.
Model SNI Task New Task
Qwen-72B-Chat 0.588 0.573
Qwen-14B-Chat 0.573 0.535
Deepseek-LLM-67B-Chat 0.529 0.504
gpt-3.5-public-turbo 0.523 0.500
Yi-34B-Chat 0.509 0.514
Table 6: Comparison of the CIF-Bench overall scores in the Public split and other leaderboards. The cells are highlighted with fading colors from maximum to minimum for the applicable numbers in a column. * indicates that the performance of pre-trained base LLMs is used to approximate the evaluation of the corresponding unavailable chat models.
Model Name CIF Public Open LLM OpenCompass
Qwen-72B-Chat 0.589 *73.60 51.90
Qwen-14B-Chat 0.564 *65.86 45.00
Deepseek-LLM-67B-Chat 0.526 71.79 42.70
gpt-3.5-Public-SFT 0.522 - 46.80
Yi-34B-Chat 0.516 65.32 47.10
Baichuan2-13B-Chat 0.512 - 32.10
Tigerbot-13B-Chat 0.494 *53.42 -
Chinese-Alpaca-2-13B 0.492 57.41 -
Chinese-Alpaca-33B 0.484 55.33 -
Ziya-Llama-13B 0.479 29.96 -
Chinese-Llama2-Linly-13B 0.479 - -
Tigerbot-7B-Chat 0.478 *47.93 -
ChatGLM3-6B 0.472 - 35.20
Chinese-Alpaca-13B 0.471 - -
ChatGLM2-6B 0.464 - -
Chinese-Alpaca-7B 0.452 48.85 -
Chinese-Alpaca-2-7B 0.448 - -
Chinese-Llama2-Linly-7B 0.443 45.44 -
Qwen-7B-Chat 0.442 *59.19 37.10
ChatGLM-6B 0.440 - -
Baichuan-13B-Chat 0.426 - -
Yi-6B-Chat 0.420 *54.08 31.90
CPM-Bee-10B 0.415 - -
Moss-Moon-003-SFT 0.399 - -
Belle-SFT-Public 0.397 - -
Telechat-7B 0.350 - -
Aquilachat-7B 0.350 - -
Baichuan2-7B-Chat 0.339 51.42 29.40
Table 7: Overall performance differences in CIF-Bench from Public to Private splits with single instructions.
Model Name Score Differencenormal-$\uparrow$ Model Name Score Differencenormal-$\uparrow$
Aquilachat-7B -0.050$\downarrow$ Chinese-Llama2-Linly-7B -0.122$\downarrow$
Baichuan-13B-Chat 0.020$\uparrow$ CPM-Bee-10B -0.178$\downarrow$
Baichuan2-13B-Chat 0.006$\uparrow$ Deepseek-LLM-67B-Chat -0.060$\downarrow$
Baichuan2-7B-Chat 0.071$\uparrow$ gpt-3.5-Public-SFT -0.187$\downarrow$
Belle-SFT-Public -0.145$\downarrow$ Moss-Moon-003-SFT -0.110$\downarrow$
ChatGLM-6B -0.112$\downarrow$ Qwen-14B-Chat -0.068$\downarrow$
ChatGLM2-6B -0.124$\downarrow$ Qwen-72B-Chat -0.068$\downarrow$
ChatGLM3-6B -0.038$\downarrow$ Qwen-7B-Chat -0.145$\downarrow$
Chinese-Alpaca-13B -0.148$\downarrow$ Telechat-7B -0.029$\downarrow$
Chinese-Alpaca-2-13B -0.171$\downarrow$ Tigerbot-13B-Chat -0.180$\downarrow$
Chinese-Alpaca-2-7B -0.138$\downarrow$ Tigerbot-7B-Chat -0.163$\downarrow$
Chinese-Alpaca-33B -0.170$\downarrow$ Yi-34B-Chat -0.014$\downarrow$
Chinese-Alpaca-7B -0.125$\downarrow$ Yi-6B-Chat -0.008$\downarrow$
Chinese-Llama2-Linly-13B -0.147$\downarrow$ Ziya-Llama-13B -0.167$\downarrow$
[…]
Aquilachat-7B -0.047$\downarrow$ -0.034$\downarrow$ Chinese-Llama2-Linly-7B -0.134$\downarrow$ -0.047$\downarrow$
Baichuan-13B-Chat 0.023$\uparrow$ 0.027$\uparrow$ CPM-Bee-10B -0.176$\downarrow$ 0.046$\uparrow$
Baichuan2-13B-Chat -0.003$\downarrow$ 0.008$\uparrow$ Deepseek-LLM-67B-Chat -0.076$\downarrow$ -0.029$\downarrow$
Baichuan2-7B-Chat 0.072$\uparrow$ 0.077$\uparrow$ gpt-3.5-Public-SFT -0.202$\downarrow$ -0.029$\downarrow$
Belle-SFT-Public -0.167$\downarrow$ -0.054$\downarrow$ Moss-Moon-003-SFT -0.124$\downarrow$ -0.028$\downarrow$
ChatGLM-6B -0.120$\downarrow$ -0.033$\downarrow$ Qwen-14B-Chat -0.088$\downarrow$ -0.038$\downarrow$
ChatGLM2-6B -0.131$\downarrow$ -0.005$\downarrow$ Qwen-72B-Chat -0.082$\downarrow$ -0.021$\downarrow$
ChatGLM3-6B -0.060$\downarrow$ -0.052$\downarrow$ Qwen-7B-Chat -0.157$\downarrow$ 0.005$\uparrow$
Chinese-Alpaca-13B -0.164$\downarrow$ -0.081$\downarrow$ Telechat-7B -0.050$\downarrow$ -0.045$\downarrow$
Chinese-Alpaca-2-13B -0.179$\downarrow$ -0.067$\downarrow$ Tigerbot-13B-Chat -0.187$\downarrow$ -0.017$\downarrow$
Chinese-Alpaca-2-7B -0.152$\downarrow$ -0.051$\downarrow$ Tigerbot-7B-Chat -0.162$\downarrow$ -0.004$\downarrow$
Chinese-Alpaca-33B -0.187$\downarrow$ -0.072$\downarrow$ Yi-34B-Chat -0.025$\downarrow$ -0.002$\downarrow$
Chinese-Alpaca-7B -0.147$\downarrow$ -0.072$\downarrow$ Yi-6B-Chat -0.022$\downarrow$ -0.012$\downarrow$
Chinese-Llama2-Linly-13B -0.153$\downarrow$ -0.031$\downarrow$ Ziya-Llama-13B -0.181$\downarrow$ -0.045$\downarrow$
5 Results Analysis
Broadly speaking, we aim to investigate the performance capabilities of current representative Chinese LLMs in a diverse set of NLP tasks to ascertain how well the annotated data with human standards with the provided instruction-following benchmark. Specifically, we ask: (i) Is our benchmark challenging enough? What kind of tasks are difficult? (ii) Is it true that LLMs perform worse when language is transferred? (iii) Do we measure the instruction-following capability well, by avoiding data contamination? (iv) Do the diverse instructions help?
Is CIF-Bench Challenging?
To ensure the reliability of our benchmark, the scores in the private split with the diversified instructions are referred to as the main results for discussion, as shown in Table 3 and Table 4. Our findings reveal that although large parameter size contributes to performance (Qwen-72B-Chat, Yi-34B-Chat, and Deepseek-LLM-67B-Chat), the effective training methods are still a boost for relatively small models such as Baichuan2-13B-Chat and Qwen-14B-Chat. Given that the highest score barely reaches $52.9$ overall out of $100$ and only 4 models exceed $50.0$, we conclude that our proposed CIF is a tough benchmark for existing LLMs for question (i).
In addition, we provide finer-grained score aggregation to further analyze the challenging task categories (n.b., most bilingual LLMs perform poorly on tasks in code, summarization, and translation categories). We point out that parallel data of Chinese and minor languages and Chinese-commented code corpora are still a scarce resource, which might lead to their poor performance on CIF-Bench’s code and translation categories. Additionally, we assume that Chinese and English bilingual LLMs, although a major branch of multilingual LLM, do not significantly benefit LLMs’ capacity to deal with minor-language-related tasks. Part of the tasks in CIF-Bench’s summarization category are very challenging, combining counterfactual reasoning and empathy estimation (i.e., task 126 and task 132 referring to Appendix A.1). Thereby, the bilingual LLMs’ poor performance on CIF-Bench’s summarization category is understandable. Detailed category-based scores on the public split are available in Table 13 in § B for further analysis.
Language Transferability.
We select the public split to investigate LLM language transferability in instruction-following. In the CIF-Bench public split, a set of 70 tasks from SNI Wang et al. (2022b) are used as representative samples of English NLP tasks equipped with directly translated input-output pairs in Chinese. We select the top-5 performing models on the public split to show the performance comparison between SNI and our $37$ original curated Chinese tasks in Table 5. Although these models maintain instruction-following capability when encountering the translated SNI data, they generally perform worse on tasks newly created in Chinese without a corresponding “copy” in English, which yields an average score decrement of $2.2\%$.
Table 9: The difference of variance of task-level scores from single to diverse instruction sets. The variance values are scaled by a factor of $1\times 10^{-3}$.
Model Name Var. Differencenormal-$\downarrow$ Model Name Var. Differencenormal-$\downarrow$
Aquilachat-7B -3.961$\downarrow$ Chinese-Llama2-Linly-7B -4.539$\downarrow$
Baichuan-13B-Chat -7.049$\downarrow$ Cpm-Bee-10B -0.661$\downarrow$
Baichuan2-13B-Chat -3.633$\downarrow$ Deepseek-Llm-67B-Chat -2.889$\downarrow$
Baichuan2-7B-Chat -1.402$\downarrow$ Gpt-3.5-Turbo-Sft -6.369$\downarrow$
Belle-13B-Sft -0.316$\downarrow$ Moss-Moon-003-Sft -6.827$\downarrow$
Chatglm-6B-Sft -5.051$\downarrow$ Qwen-14B-Chat -0.978$\downarrow$
Chatglm2-6B -3.980$\downarrow$ Qwen-72B-Chat -1.817$\downarrow$
Chatglm3-6B -0.413$\downarrow$ Qwen-7B-Chat -3.185$\downarrow$
Chinese-Alpaca-13B -8.303$\downarrow$ Telechat-7B -6.090$\downarrow$
Chinese-Alpaca-2-13B -4.814$\downarrow$ Tigerbot-13B-Chat -3.816$\downarrow$
Chinese-Alpaca-2-7B -4.494$\downarrow$ Tigerbot-7B-Chat -6.004$\downarrow$
Chinese-Alpaca-33B -5.000$\downarrow$ Yi-34B-Chat -1.942$\downarrow$
Chinese-Alpaca-7B -2.961$\downarrow$ Yi-6B-Chat -6.397$\downarrow$
Chinese-Llama2-Linly-13B -2.961$\downarrow$ Ziya-Llama-13B -3.001$\downarrow$
Data Contamination Does Exist.
As mentioned in §3, we evaluate the model performances on the public split with half of the input-output pairs in the single instruction setting, with which we can conveniently probe the benchmark data contamination issue of the LLMs.
We first compare the CIF-Bench public results with two comprehensive LLM benchmarks, including the Open LLM Leaderboard Beeching et al. (2023), as well as an English-Chinese leaderboard, OpenCompass Contributors (2023). As suggested in Table 6 with rows ranked in the descending order of the overall public scores, the results on CIF-Bench are aligned with the other two popular benchmarks, which therefore verifies the reliability of our evaluation pipeline. However, we suspect the highly correlative rankings could be a result of the benchmark data leakage in those “web-scale” pre-training data, since $117$ of the constructed tasks and instances in the public split are sourced from the internet.
To further confirm such suspicions, we calculate the performance changes of overall scores in the same single instruction setting, but with different input-output pairs from the public and private splits. Revealed by the differences in Table 7, there is a noticeable performance drop for most (25/28) of the models when a large part of the data translated from public sources is replaced by our original annotations. Consequently, incoming models submitted to the proposed CIF-Bench will restricted to the private split for the sake of evaluation reliability.
It is likely that both the leakage of the input-output instances and the tasks themselves contribute to the mentioned evaluation bias. To compare the two factors for the downgraded performances, we analyze the performance shift with the $113$ “Existing” tasks translated from English or originally in Chinese and the $37$ “New” tasks we crafted from scratch. As revealed in Table 8, the LLMs have impaired performance when given newly curated data instances for a set of seen “Existing” tasks, yielding an average $11.0\%$ score decrease. In contrast, these models on average perform $2.5\%$ worse, with both definitely-unseen tasks and corresponding input-output pairs. We hence conclude that the leakage of the data instances plays a more significant role than the tasks themselves in evaluation biases.
Instruction Diversity for Evaluation Robustness.
With the motivation that a model might produce inconsistent output given various instruction,input holding the same semantics, we argue that a diversified instruction set can increase the evaluation robustness by incorporating more corner cases. We separately calculate the task-level score variance in the private split for the conditions of using one and five instructions to verify the improvement. We find that increasing the diversity of the task instructions can bring extra robustness to the evaluation, as the evaluation scores are stabilized to lower variance for all the tested LLMs (see in Table 9).
Human Annotation for Verification.
To verify the annotation quality and reliability , we invite 3 annotators with expert-level NLP research backgrounds to assess the model outputs in public split with the same task-level instruction. The evaluation dimensions include: “Faithfulness”: human experts reflect on the absolute quality of a model’s output in a binary (yes/no) form. “Level of preference”: a 5-point Likert scale was provided to the experts to assess the relative quality of the model outputs. We randomly sample tasks according to the task category distribution, and pick three models performing differently in general, specifically Moss-Moon-003-sft ($0.399$), Baichuan-13B-Chat ($0.426$), and Qwen-72B-Chat ($0.589$). Considering the diverse and open-ended task, we first measure quality by comparing the pairwise agreement between each two annotators, reporting an average agreement of 0.49. Furthermore, we employ Cohen’s kappa (Ben-David, 2008) to measure inter-rater reliability, reporting an average of 0.3729 across the $153$ questions, implying that the results are substantially reliable. Specifically, the experts scored $0.4966$ on the dichotomous form yet $0.2492$ on the more varied options, suggesting that completing $153$ questions is challenging even for human experts. We further explore the correlation between the model prediction with human evaluation(Spearman’s $r=0.4043$), suggesting that most annotated were indeed truthful and the models can be relied upon to generate output for this task.
6 Conclusion
In summary, CIF-Bench not only exposes the limitations of current LLMs in navigating the complexities of Chinese language instruction-following tasks but also provides a foundational platform for future advancements in LLM generalizability research. Through this work, we aim to facilitate the development of more adaptable, culturally aware, and linguistically diverse language models, capable of truly understanding and interacting with the global tapestry of human language.
Limitations
Recruiting human subjects for annotation limits the reproducibility of human evaluation. In addition, we recognize that there might be more suitable baseline models, whilst in this study only a few of the most advanced models were used. Finally, despite annotation and discrimination by human experts, there may still be offensive content in the data due to both human education and environmental factors. It is worth noting, however, that identifying offensive language is not the purpose of this work.
Ethics Statement
The dataset presented was annotated by a third-party professional annotation company. During the annotation process, we considered the following aspects to ensure the protection of the annotators. (1) Consent: To ensure that our participants agreed to the annotation task, we asked them to read the task guidelines and instructions before starting the work. If they felt uncomfortable, they could withdraw from the task at any time. (2) Confidentiality: The entire annotation process was anonymous and we did not know any information about the participants in the task. (3) Assurance: all data were obtained from open-source datasets or resources.
Appendix A Task Details
A.1 Full List of Tasks and Evaluation
We provide a full list of the task names and the source for input-output annotation in this subsection. The comprehensive task descriptions and the corresponding evaluation prompts can be found in the supplementary files.
Table 10: Full task list and source (1/3).
Task ID & Name Source
0 Negotiation Strategy Detection
SNI Wang et al. (2022b)
1 Grammar Error Correction
SNI Wang et al. (2022b)
2 Overlap Extraction
SNI Wang et al. (2022b)
3 Commonsense
SNI Wang et al. (2022b)
4 Data to Text
SNI Wang et al. (2022b)
5 Keyword Tagging
SNI Wang et al. (2022b)
6 Answerability Classification
SNI Wang et al. (2022b)
7 Dialogue Act Recognition
SNI Wang et al. (2022b)
8 Cause Effect Classification
SNI Wang et al. (2022b)
9 Question Rewriting
SNI Wang et al. (2022b)
10 Textual Entailment
SNI Wang et al. (2022b)
11 Coreference Resolution
SNI Wang et al. (2022b)
12 Title Generation
SNI Wang et al. (2022b)
13 Entity Relation Classification
SNI Wang et al. (2022b)
14 Punctuation Error Detection
SNI Wang et al. (2022b)
15 Style Transfer
SNI Wang et al. (2022b)
16 Sentence Expansion
SNI Wang et al. (2022b)
17 Poem Generation
SNI Wang et al. (2022b)
18 Discourse Relation Classification
SNI Wang et al. (2022b)
19 Mathematics
SNI Wang et al. (2022b)
20 Text Simplification
SNI Wang et al. (2022b)
21 Sentence Compression
SNI Wang et al. (2022b)
22 Spelling Error Detection
SNI Wang et al. (2022b)
23 Irony Detection
SNI Wang et al. (2022b)
24 Number Conversion
SNI Wang et al. (2022b)
25 Word Relation Classification
SNI Wang et al. (2022b)
26 Paraphrasing
SNI Wang et al. (2022b)
27 Grammar Error Detection
SNI Wang et al. (2022b)
28 Text Matching
SNI Wang et al. (2022b)
29 Fill in The Blank
SNI Wang et al. (2022b)
30 Speaker Relation Classification
SNI Wang et al. (2022b)
31 Entity Generation
SNI Wang et al. (2022b)
32 Summarization
SNI Wang et al. (2022b)
33 Spam Classification
SNI Wang et al. (2022b)
34 Stereotype Detection
SNI Wang et al. (2022b)
35 Dialogue State Tracking
SNI Wang et al. (2022b)
36 Dialogue State Tracking
SNI Wang et al. (2022b)
37 Sentence Perturbation
SNI Wang et al. (2022b)
38 Text Quality Evaluation
SNI Wang et al. (2022b)
39 Linguistic Probing
SNI Wang et al. (2022b)
40 Information Extraction
SNI Wang et al. (2022b)
41 Emotion Prediction
SNI Wang et al. (2022b)
42 Discourse Connective Identification
SNI Wang et al. (2022b)
43 Question Generation
SNI Wang et al. (2022b)
44 Stance Detection
SNI Wang et al. (2022b)
45 Sentiment Analysis
SNI Wang et al. (2022b)
46 Story Composition
SNI Wang et al. (2022b)
47 Program Execution
SNI Wang et al. (2022b)
48 Gender Classification
SNI Wang et al. (2022b)
49 Named Entity Recognition
SNI Wang et al. (2022b)
50 Toxic Language Detection
SNI Wang et al. (2022b)
51 Question Decomposition
SNI Wang et al. (2022b)
52 Sentence Ordering
SNI Wang et al. (2022b)
53 Text to Code
SNI Wang et al. (2022b)
54 Fact Verification
SNI Wang et al. (2022b)
55 Speaker Identification
SNI Wang et al. (2022b)
56 Answer Verification
SNI Wang et al. (2022b)
57 Wrong Candidate Generation
SNI Wang et al. (2022b)
58 Dialogue Generation
SNI Wang et al. (2022b)
59 Text Completion
SNI Wang et al. (2022b)
60 Pos Tagging
SNI Wang et al. (2022b)
Table 11: Full task list and source (2/3).
Task ID & Name Source
61 Explanation
SNI Wang et al. (2022b)
62 Sentence Composition
SNI Wang et al. (2022b)
63 Question Understanding
SNI Wang et al. (2022b)
64 Intent Identification
SNI Wang et al. (2022b)
65 Word Semantics
SNI Wang et al. (2022b)
66 Code to Text
SNI Wang et al. (2022b)
67 Preposition Prediction
SNI Wang et al. (2022b)
68 Text Categorization
SNI Wang et al. (2022b)
69 Question Answering
SNI Wang et al. (2022b)
70 Commonsense Classification
N/A
71 Ancient Chinese Poem Retrieval
N/A
72 Ancient Chinese Translation
N/A
73 Chinese Rhyme Detection
N/A
74 Nationality Detection
N/A
75 Region Detection
N/A
76 Chinese Idiom Explanation
N/A
77 Name Allusion Detection
N/A
78 Chinese Ambiguity Sentence Location
N/A
79 Chinese Winograd Schema Challenge
FewCLUE Xu et al. (2021)
80 Chinese Modern Abbreviation Explanation
Wikipedia Wikipedia (2024)
81 Chinese Epigraph Detection
N/A
82 Chinese Dialect Translation
N/A
83 Chinese Attractions List
N/A
85 Chinese Typo Categorization
N/A
86 Chinese Fiction Characteristic Detection
N/A
87 Chinese Figurative Detection
N/A
88 Chinese Metaphor Explanation
N/A
89 Chinese Medicine Detection
N/A
90 Chinese Pinyin Detection
N/A
91 Chinese Wubi Written
N/A
92 Intimacy Score Prediction
Pei and Jurgens (2020)
93 Sentence Level Uncertainty Judgement
Pei and Jurgens (2021)
94 Chinese Relative Identification
N/A
96 Chinese Heteronomous Language Detection
N/A
97 Code Debug
https://blog.csdn.net
98 Code Translate
https://leetcode.cn
99 Function Explanation
https://www.liaoxuefeng.com/
100 Bias Detoxication
CORGI-PM Zhang et al. (2023a)
101 MultiLabel Chinese Humor Categorization
Tseng et al. (2020)
102 Legal Term Retrieval
N/A
103 Patronizing Condescending Multilabel
Perez-Almendros et al. (2022)
104 CommonSense Explanation
Wang et al. (2020)
105 Event Type Detection
MUSIED (Xi et al., 2022)
106 Argument Mining
N/A
107 Theory of Mind
Big-Bench Theory of Mind (bench authors, 2023)
108 Game Playing
Big-Bench Language Games (bench authors, 2023)
109 IQ Test
Huang et al. (2023a)
110 Joke Explanation
N/A
111 Role Playing
TRPG https://bilibili.com Côté et al. (2018)
112 Text De-Identification
N/A
113 Outline Generation
N/A
114 Pros Cons Listing
N/A
115 Joke Telling
N/A
116 Affordance
N/A
Table 12: Full task list and source (2/3).
Task ID & Name Source
117 Material Synthesis
Bara et al. (2021)
118 Tool use
Taskmatrix Liang et al. (2023)
119 Concept Abstraction
N/A
120 Rhyme Aligned Generation
N/A
121 Advertising
N/A
122 Mind Tree Generation
N/A
123 First Order Logic
FOLIO Han et al. (2022)
124 Critical Thinking
N/A
125 Empathy Detection
N/A
126 Social Norms Detection
Moral Stories Emelin et al. (2021)
127 Make Positive
Ziems et al. (2022)
128 Translate to Ancient Chinese
N/A
129 Recipe Generation
Market (2018)
130 Imagination
N/A
131 Compositional Reasoning
Lake and Baroni (2017)
132 Personality Detection
N/A
133 Table Generation
N/A
134 Flowchart Generation
N/A
135 Review Generation
N/A
136 Draw Figure with symbol
N/A
137 CommonsenseQA
CommonsenseQA Talmor et al. (2018)
138 ReadingComprehensionQA
Rajpurkar et al. (2018)
139 DiscreteOperationQA
DROPDua et al. (2019)
140 MultiHopQA
HotpotQA Yang et al. (2018)
141 CommonsenseNLI
HellaSwag Zellers et al. (2019)
142 ConversationalQA
CoQA Reddy et al. (2019)
143 MathQA
GSM8K Cobbe et al. (2021)
144 English translation
N/A
145 French translation
DiaBLa Bawden et al. (2021)
146 Arabic translation
Altammami et al. (2020)
147 Japanese translation
Oda (2016)
148 Spanish translation
"European Commission and Technology." (2017)
149 Bengali translation
Islam et al. (2022)
150 Tamil translation
Ramasamy et al. (2012)
151 Gujarati translation
Shah and Bakrola (2019)
A.2 Category Description
We provide the task category description in this subsection.
Chinese Culture (18).
Focuses on aspects unique to Chinese history, society, and language, therefore testing the model’s understanding of cultural nuances.
Classification (21).
Addresses classification tasks, such as determining correctness or whether something belongs to a specific category.
Code (5).
Tests the model’s proficiency in understanding and generating computer code across various programming languages.
Commonsense (36).
Evaluates the model’s grasp of general knowledge and everyday reasoning that humans consider obvious.
Creative Natural Language Generation (NLG) (21).
Measures the model’s ability to produce imaginative and novel text outputs, ranging from stories to creative descriptions.
Evaluation (5).
Focuses on assessing other models or systems, therefore testing the ability to judge and provide feedback on performance.
Grammar (10).
Assesses the model’s understanding of linguistic rules and its ability to apply them correctly in text generation.
Linguistic (24).
Involves tasks that test the model’s understanding of language structure, including syntax, semantics, and morphology.
Motion Detection (6).
Uncommon in LLMs, this refers to tasks related to interpreting descriptions of motion or predicting outcomes based on textual motion descriptions.
Named Entity Recognition (NER) (12).
Involves identifying and categorizing key information (e.g., names, places, dates) within the text.
Natural Language Inference (NLI) (30).
Tests the model’s ability to understand relationships between sentences, such as contradiction, entailment, and neutrality.
Question Answering (QA) (19).
Evaluates the model’s ability to understand and respond to questions with accurate and relevant answers.
Reasoning (29).
Involves tasks that require logical thinking, problem-solving, and deduction to arrive at correct conclusions.
Role Playing (2).
Tests the model’s ability to adopt personas or roles in conversational contexts, assessing its versatility in generating context-appropriate responses.
Sentiment (8).
Evaluates the model’s ability to detect and interpret emotional tones in text, such as positive, negative, or neutral sentiments.
Structured Data (16).
Involves interpreting and generating responses based on structured information such as tables, charts, and databases.
Style Transfer (20).
Tests the model’s ability to convert text from one stylistic or tonal form to another while retaining the original content’s meaning.
Summarization (9).
Assesses the model’s ability to condense longer texts into shorter, coherent summaries capturing the essential points.
Toxicity (3).
Focuses on identifying and mitigating harmful or stereotypical content in text generation.
Translation (9).
Evaluates the model’s ability to accurately translate text between languages, testing its linguistic versatility and understanding.
Appendix B CIF-Bench Results in Public Split
We provide the category-based results in public split here.
Table 13: Overall Results in CIF-Bench Public Split with Single Instruction. The first column is the average score across all the tasks, and the rest columns are average scores grouped by task categories. The cells are highlighted with fading colors from maximum to minimum in a column.
Model Name Overall Chinese Culture Classification Code Commonsense Creative NLG Evaluation Grammar Linguistic Motion Detection NER
Qwen-72B-Chat 0.589 0.512 0.716 0.444 0.706 0.587 0.661 0.424 0.521 0.694 0.515
Qwen-14B-Chat 0.564 0.481 0.678 0.416 0.657 0.567 0.669 0.396 0.485 0.663 0.486
Deepseek-LLM-67B-Chat 0.526 0.477 0.617 0.364 0.609 0.559 0.573 0.374 0.458 0.631 0.493
GPT-3.5-Public-SFT 0.522 0.316 0.611 0.492 0.578 0.538 0.639 0.377 0.447 0.580 0.492
Yi-34B-Chat 0.516 0.452 0.607 0.437 0.624 0.516 0.545 0.254 0.382 0.671 0.398
Baichuan2-13B-Chat 0.512 0.446 0.623 0.403 0.600 0.505 0.582 0.352 0.423 0.633 0.435
Tigerbot-13B-Chat 0.494 0.350 0.558 0.447 0.599 0.528 0.707 0.352 0.447 0.551 0.498
Chinese-Alpaca-2-13B 0.492 0.260 0.572 0.434 0.533 0.562 0.574 0.318 0.417 0.624 0.467
Chinese-Alpaca-33B 0.484 0.274 0.546 0.470 0.527 0.540 0.703 0.332 0.382 0.582 0.464
Ziya-Llama-13B 0.479 0.287 0.550 0.422 0.523 0.551 0.650 0.294 0.384 0.610 0.437
Chinese-Llama2-Linly-13B 0.479 0.286 0.623 0.439 0.549 0.535 0.626 0.286 0.403 0.587 0.468
Tigerbot-7B-Chat 0.478 0.354 0.528 0.440 0.570 0.540 0.708 0.314 0.430 0.528 0.413
ChatGLM3-6B 0.472 0.321 0.488 0.436 0.527 0.503 0.588 0.290 0.328 0.574 0.415
Chinese-Alpaca-13B 0.471 0.264 0.553 0.443 0.495 0.525 0.587 0.334 0.394 0.653 0.457
ChatGLM2-6B 0.464 0.334 0.532 0.436 0.522 0.527 0.651 0.314 0.395 0.536 0.402
Chinese-Alpaca-7B 0.452 0.237 0.536 0.438 0.484 0.502 0.672 0.318 0.389 0.652 0.394
Chinese-Alpaca-2-7B 0.448 0.251 0.472 0.435 0.480 0.532 0.577 0.268 0.348 0.596 0.431
Chinese-Llama2-Linly-7B 0.443 0.264 0.558 0.419 0.497 0.522 0.664 0.236 0.381 0.593 0.381
Qwen-7B-Chat 0.442 0.313 0.549 0.404 0.520 0.515 0.646 0.244 0.411 0.570 0.368
ChatGLM-6B 0.440 0.311 0.499 0.446 0.484 0.548 0.558 0.278 0.382 0.484 0.386
Baichuan-13B-Chat 0.426 0.355 0.416 0.361 0.516 0.416 0.564 0.324 0.374 0.380 0.394
Yi-6B-Chat 0.420 0.320 0.439 0.395 0.489 0.449 0.493 0.230 0.293 0.587 0.341
CPM-Bee-10B 0.415 0.382 0.455 0.284 0.431 0.508 0.300 0.317 0.367 0.494 0.397
Moss-Moon-003-SFT 0.399 0.233 0.465 0.389 0.427 0.482 0.509 0.274 0.369 0.526 0.385
Belle-SFT-Public 0.397 0.196 0.503 0.376 0.426 0.472 0.543 0.269 0.371 0.512 0.356
Telechat-7B 0.350 0.172 0.299 0.438 0.386 0.456 0.400 0.138 0.202 0.412 0.322
Aquilachat-7B 0.350 0.203 0.270 0.357 0.404 0.449 0.394 0.090 0.260 0.348 0.322
Baichuan2-7B-Chat 0.339 0.345 0.595 0.154 0.455 0.327 0.523 0.362 0.354 0.466 0.233
Model Name Overall NLI QA Reasoning Role Playing Sentiment Structured Data Style Transfer Summarization Toxic Translation
Qwen-72B-Chat 0.589 0.695 0.668 0.539 0.752 0.637 0.505 0.587 0.609 0.671 0.466
Qwen-14B-Chat 0.564 0.647 0.609 0.498 0.757 0.638 0.460 0.610 0.629 0.691 0.467
Deepseek-LLM-67B-Chat 0.526 0.588 0.624 0.444 0.694 0.592 0.384 0.576 0.594 0.666 0.439
GPT-3.5-Public-SFT 0.522 0.587 0.565 0.498 0.745 0.583 0.444 0.501 0.620 0.643 0.452
Yi-34B-Chat 0.516 0.631 0.592 0.460 0.761 0.566 0.440 0.551 0.610 0.608 0.408
Baichuan2-13B-Chat 0.512 0.600 0.591 0.474 0.751 0.597 0.434 0.525 0.572 0.494 0.372
Tigerbot-13B-Chat 0.494 0.571 0.569 0.413 0.732 0.560 0.365 0.502 0.607 0.601 0.306
Chinese-Alpaca-2-13B 0.492 0.566 0.545 0.420 0.712 0.595 0.382 0.488 0.641 0.740 0.347
Chinese-Alpaca-33B 0.484 0.550 0.506 0.423 0.732 0.548 0.342 0.494 0.629 0.648 0.334
Ziya-Llama-13B 0.479 0.546 0.499 0.404 0.749 0.582 0.367 0.499 0.629 0.722 0.313
Chinese-Llama2-Linly-13B 0.479 0.563 0.524 0.411 0.676 0.561 0.359 0.482 0.602 0.696 0.313
Tigerbot-7B-Chat 0.478 0.532 0.554 0.393 0.731 0.583 0.351 0.519 0.630 0.614 0.291
ChatGLM3-6B 0.472 0.557 0.526 0.397 0.749 0.612 0.431 0.529 0.620 0.589 0.392
Chinese-Alpaca-13B 0.471 0.524 0.513 0.402 0.726 0.526 0.323 0.486 0.628 0.702 0.336
ChatGLM2-6B 0.464 0.520 0.533 0.407 0.725 0.506 0.363 0.480 0.627 0.661 0.303
Chinese-Alpaca-7B 0.452 0.504 0.501 0.351 0.699 0.543 0.365 0.478 0.623 0.711 0.328
Chinese-Alpaca-2-7B 0.448 0.509 0.493 0.344 0.703 0.510 0.334 0.483 0.637 0.596 0.343
Chinese-Llama2-Linly-7B 0.443 0.496 0.546 0.350 0.713 0.559 0.323 0.495 0.603 0.584 0.293
Qwen-7B-Chat 0.442 0.489 0.514 0.384 0.713 0.563 0.328 0.463 0.576 0.639 0.281
ChatGLM-6B 0.440 0.480 0.483 0.353 0.738 0.460 0.346 0.480 0.633 0.543 0.322
Baichuan-13B-Chat 0.426 0.531 0.584 0.339 0.668 0.478 0.402 0.459 0.559 0.497 0.392
Yi-6B-Chat 0.420 0.496 0.516 0.344 0.742 0.488 0.348 0.498 0.627 0.510 0.285
CPM-Bee-10B 0.415 0.451 0.472 0.304 0.647 0.329 0.284 0.538 0.534 0.486 0.305
Moss-Moon-003-SFT 0.399 0.403 0.457 0.325 0.712 0.450 0.304 0.435 0.594 0.542 0.308
Belle-SFT-Public 0.397 0.450 0.430 0.338 0.645 0.426 0.300 0.398 0.558 0.683 0.224
Telechat-7B 0.350 0.375 0.414 0.261 0.660 0.341 0.320 0.462 0.639 0.494 0.304
Aquilachat-7B 0.350 0.385 0.426 0.274 0.595 0.308 0.267 0.434 0.607 0.409 0.355
Baichuan2-7B-Chat 0.339 0.414 0.349 0.339 0.673 0.429 0.300 0.246 0.097 0.357 0.130